# FastDrag: Manipulate Anything in One Step

### v2

@@ 1 Introduction
However, existing methods typically involve $n$-step iterative semantic optimization in latent space to obtain optimized latent with desired semantic based on the user-provided drag instructions. These optimization approaches can be categorized primarily into motion-based [22, 27, 14, 15, 31, 4] and gradient-based [19, 18] $n$-step iterative optimizations, as shown in Fig. 1(a). $n$-step iterations in motion-based methods are necessary to avoid abrupt changes in the latent space, preventing image distortions and ensuring a stable optimization process. This is exemplified in DragDiffusion [27] and GoodDrag [31], which require 70 to 80 iterations of point tracking and motion supervision for optimization. In addition, gradient-based methods align sampling results with the drag instructions through gradient guidance [6]. In this way, they also require multiple steps due to the optimizer [12, 24] needing multiple iterations for non-convex optimization. For instance, DragonDiffusion [19] requires around 50 gradient steps to accomplish the latent optimization. Therefore, existing drag-based image editing methods often suffer from significant time consumption due to $n$-step iterations required for latent semantic optimization, thus limiting the practical applications.
@@
Figure 1: (a) Existing methods usually require multiple iterations to transform an image from its original semantic to desired semantic; (b) Our method utilizes latent warpage function (LWF) to calculate the warpage vectors (i.e., $\bm{v}_{j}$) to move each individual pixel on feature map and achieve semantic optimization in a single step.
@@ 2.2 Drag-based Image Editing
Drag-based image editing achieves precise spatial control over specific regions of the image based on user-provided drag instructions. Existing drag-based image editing methods generally rely on $n$-step latent semantic optimization in latent space to achieve image editing. These methods fall into two main categories: motion-based [22, 27, 31, 4, 15, 14] and gradient-based [19, 18] optimizations. For example, DragGAN [22] employs generative adversarial network (GAN) for drag-based image editing with iterative point tracking and motion supervision steps. However, the image quality of the methods using GAN for image generation is worse than diffusion models [5]. Therefore, a series of diffusion-based methods have been proposed for drag-based image editing. For instance, DragDiffusion [27] employs iterative point tracking and motion supervision for latent semantic optimization to achieve drag-based editing. Building on this foundation, GoodDrag [31], StableDrag [4], DragNoise [15], and FreeDrag [14] have made significant improvements to the motion-based methods. Without coincidence, by utilizing feature correspondences, DragonDiffusion [19] and its improved version DiffEditor [18] formulate an energy function that conforms to the desired editing results, thereby transforming the image editing task into a gradient-based process that enables drag-based editing. However, these methods inherently require $n$-step iterations for latent optimization, which significantly increase the time consumption. Although SDEDrag [21] does not require $n$-step iterative optimization, it is still time-consuming due to the stochastic differential equation (SDE) process for diffusion. To this end, based on latent diffusion model (LDM) [25], we propose a novel one-step optimization method that substantially accelerates the image editing speeds.
@@ 3.2.1 Warpage Vector Calculation using LWF
In drag-based image editing, each drag instruction $\bm{d}_{i}$ in a set of $k$ drag instructions $\bm{D}=\{\bm{d}_{i}\mid i=1,\ldots,k;k\in\mathbb{Z}\}$ can simultaneously influence a feature point $p_{j}$ on the mask region $\bm{P}=\{p_{j}\mid j=1,\ldots,m;m\in\mathbb{Z}\}$ provided by the user. To get a uniquely determined vector, i.e., warpage vector $\bm{{v}_{j}}$ to adjust the position of feature point $p_{j}$ (will be discussed in Sec.3.2.2), we propose a latent warpage function $f_{LWF}(\cdot)$ to aggregate multiple component warpage vectors caused by different drag instructions, i.e., $\bm{v}_{j}^{i*}$, with balanced weights to avoid deviating from the desired drag effect. The function is given as follows:
@@
$f_{LWF}(\bm{P},\bm{D},j)=\sum_{i}^{k}{w_{j}^{i}\cdot\lambda_{j}^{i}\cdot\bm{d}_{i}},$ (7)
@@ 3.3 Bilateral Nearest Neighbor Interpolation
To enhance the semantic integrity, BNNI interpolates points in null region using similar features from their neighboring areas, thus ensuring the semantic integrity and enhancing the quality of drag editing. Let $\mathcal{N}$ be a point with coordinate $(x_{\mathcal{N}},y_{\mathcal{N}})$ in null regions, we identify the nearest points of $\mathcal{N}$ containing value in four directions: up, right, down, and left, as illustrated in Fig. 5, which are used as reference points for interpolation. Then, the interpolated value for null point $\mathcal{N}$ can be calculated as:
@@ 4.1 Qualitative Evaluation
Additionally, we perform multi-point dragging experiments, illustrated in the fourth row of Fig. 6. Both DragDiffusion and DragNoise fail to stretch the back of the sofa, while FreeDrag incorrectly stretches unintended parts of the sofa. Through the LWF introduced in Sec. 3.2.1, FastDrag can manipulate all dragged points to their target locations while preserving the content in unmasked regions. More results of FastDrag are illustrated in supplementary Sec. C.
@@ 4.2 Quantitative Comparison
To better demonstrate the superiority of FastDrag, we conduct quantitative comparison using DragBench dataset, which consists of 211 different types of images with 349 pairs of handle and target points. Here, mean distance (MD) [22] and image fidelity (IF) [10] are employed as performance metrics, where MD evaluates the precision of drag editing, and IF measures the consistency between the generated and original images by averaging the learned perceptual image patch similarity (LPIPS) [30]. Specifically, 1-LPIPS is employed as the IF metric in our experiment to facilitate comparison. In addition, we compare the average time required per point to demonstrate the time efficiency of our proposed FastDrag. The results are given in Fig. 8.
@@
Apart from [27],[15],[14], two other state-of-the-art methods, i.e., GoodDrag [31] and DiffEditor [18], are also adopted for comparison, with DiffEditor being the current fastest drag-based editing method. Due to well-designed one-step warpage optimization and consistency-preserving strategy, our FastDrag does not require LoRA training preparation, resulting in significantly reduced time consumption (i.e., 3.12 seconds), which is nearly 700% faster than DiffEditor (i.e., 21.68 seconds), and 2800% faster than the typical baseline DragDiffusion (i.e., 1 min and 21.54 seconds). Moreover, even using standard U-Net without LCM, our method is still much faster than DiffEditor and far outperforms all other state-of-the-art methods. It is particularly noteworthy that, even with an A100 GPU, DiffEditor still requires 13.88 seconds according to [18], whereas our FastDrag is evaluated with an RTX 3090.
@@ 4.3 Ablation Study
Inversion Step: To determine the number of inversion steps in diffusion inversion with LCM-equipped U-Net, we conduct an ablation experiment with number of inversion steps set as $t$ = 4, 6, 8, 10, 12, 14, 20, and 30, where IF and MD are used to evaluate the balance between the consistency with original image and the desired drag effects. The results are given in Fig. 9 and Fig. 8, where we can see that when $t<6$, the generated images lack sufficient detail to accurately reconstruct the original images. Conversely, when $t>6$, it can successfully recover complex details such as intricate fur textures while maintaining high image quality. However, when $t>14$, the images become excessively similar to the original, which negatively impacts the effectiveness of the drag effect. By comprehensive evaluation of both the drag effect and the similarity to the original images, we select 10 as the number of inversion steps for our method with LCM to balance the drag effect.
@@ Appendix A Supplementary Experiments
Appendix B Implementation Details
We utilize a pretrained LDM (i.e., Stable Diffusion 1.5 [26]) as our diffusion model, where the U-Net structure is adapted with LCM-distilled weights from Stable Diffusion 1.5. Unless otherwise specified, the default setting for inversion and sampling step is 10. Following DragDiffusion [27], classifier-free guidance (CFG) [8] is not applied in diffusion model, and we optimize the diffusion latent at the 7th step. All other configurations follow that used in DragDiffusion. Our experiments are conducted on an RTX 3090 GPU with 24G memory.
Appendix C More Results

### v3

@@ 1 Introduction
However, existing methods typically involve $n$-step iterative semantic optimization in latent space to obtain optimized latent with desired semantic based on the user-provided drag instructions. They focus on optimizing a small region of the image at each step, requiring $n$ small-scale and short-distance adjustments to achieve overall latent optimization, leading to a significant amount of time. These optimization approaches can be categorized primarily into motion-based [23, 28, 15, 16, 32, 4] and gradient-based [20, 19] $n$-step iterative optimizations, as shown in Fig. 1(a). $n$-step iterations in motion-based methods are necessary to avoid abrupt changes in the latent space, preventing image distortions and ensuring a stable optimization process. This is exemplified in DragDiffusion [28] and GoodDrag [32], which require 70 to 80 iterations of point tracking and motion supervision for optimization. In addition, gradient-based methods align sampling results with the drag instructions through gradient guidance [6]. In this way, they also require multiple steps due to the optimizer [13, 25] needing multiple iterations for non-convex optimization. For instance, DragonDiffusion [20] requires around 50 gradient steps to accomplish the latent optimization. Therefore, existing drag-based image editing methods often suffer from significant time consumption due to $n$-step iterations required for latent semantic optimization, thus limiting the practical applications.
@@
Figure 1: (a) Existing methods usually require multiple iterations to transform an image from its original semantic to desired semantic; (b) Our method utilizes latent warpage function (LWF) to calculate the warpage vectors (i.e., $\bm{v}_{j}$) to move each individual pixel on feature map and achieve semantic optimization in one step.
@@ 2.2 Drag-based Image Editing
Drag-based image editing achieves precise spatial control over specific regions of the image based on user-provided drag instructions. Existing drag-based image editing methods generally rely on $n$-step latent semantic optimization in latent space to achieve image editing. These methods fall into two main categories: motion-based [23, 28, 32, 4, 16, 15, 9] and gradient-based [20, 19] optimizations. For example, DragGAN [23] employs generative adversarial network (GAN) for drag-based image editing with iterative point tracking and motion supervision steps. However, the image quality of the methods using GAN for image generation is worse than diffusion models [5]. Therefore, a series of diffusion-based methods have been proposed for drag-based image editing. For instance, DragDiffusion [28] employs iterative point tracking and motion supervision for latent semantic optimization to achieve drag-based editing. Building on this foundation, GoodDrag [32], StableDrag [4], DragNoise [16], and FreeDrag [15] have made significant improvements to the motion-based methods. Without coincidence, by utilizing feature correspondences, DragonDiffusion [20] and its improved version DiffEditor [19] formulate an energy function that conforms to the desired editing results, thereby transforming the image editing task into a gradient-based process that enables drag-based editing. However, these methods inherently require $n$-step iterations for latent optimization, which significantly increases the time consumption. Although SDEDrag [22] does not require $n$-step iterative optimization, it is still time-consuming due to the stochastic differential equation (SDE) process for diffusion. In addition, while EasyDrag [9] offers user-friendship editing, its requirement for over 24GB of memory (i.e., a 3090 GPU) limits its broad applicability. To this end, based on latent diffusion model (LDM) [26], we propose a novel one-step optimization method that substantially accelerates the image editing speeds.
@@ 3.2.1 Warpage Vector Calculation using LWF
In drag-based image editing, each drag instruction $\bm{d}_{i}$ in a set of $k$ drag instructions $\bm{D}=\{\bm{d}_{i}\mid i=1,\ldots,k;k\in\mathbb{Z}\}$ can simultaneously influence a feature point $p_{j}$ on the mask region $\bm{P}=\{p_{j}\mid j=1,\ldots,m;m\in\mathbb{Z}\}$ provided by the user. As shown in Fig. 3, the mask region is represented by the brighter areas in the image, indicating the specific image area to be edited. To get a uniquely determined vector, i.e., warpage vector $\bm{{v}_{j}}$ to adjust the position of feature point $p_{j}$ (will be discussed in Sec.3.2.2), we propose a latent warpage function $f_{LWF}(\cdot)$ to aggregate multiple component warpage vectors caused by different drag instructions, i.e., $\bm{v}_{j}^{i*}$, with balanced weights to avoid deviating from the desired drag effect. The function is given as follows:
@@
$f_{LWF}(\bm{P},\bm{D},j)=\sum_{i}^{k}{w_{j}^{i}\cdot\lambda_{j}^{i}\cdot\bm{d}_{i}}$ (7)
@@ 3.3 Bilateral Nearest Neighbor Interpolation
To enhance the semantic integrity, BNNI interpolates points in null region using similar features from their neighboring areas in horizontal and vertical directions, thus ensuring the semantic integrity and enhancing the quality of drag editing. Let $\mathcal{N}$ be a point with coordinate $(x_{\mathcal{N}},y_{\mathcal{N}})$ in null regions, we identify the nearest points of $\mathcal{N}$ containing value in four directions: up, right, down, and left, as illustrated in Fig. 5, which are used as reference points for interpolation. Then, the interpolated value for null point $\mathcal{N}$ can be calculated as:
@@ 4.1 Qualitative Evaluation
Additionally, we perform multi-point dragging experiments, illustrated in the fourth row of Fig. 6. Both DragDiffusion and DragNoise fail to stretch the back of the sofa, while FreeDrag incorrectly stretches unintended parts of the sofa. Through the LWF introduced in Sec. 3.2.1, FastDrag can manipulate all dragged points to their target locations while preserving the content in unmasked regions. More results of FastDrag are illustrated in supplementary Sec. E.
@@ 4.2 Quantitative Comparison
To better demonstrate the superiority of FastDrag, we conduct quantitative comparison using DragBench dataset [28], which consists of 205 different types of images with 349 pairs of handle and target points. Here, mean distance (MD) [23] and image fidelity (IF) [11] are employed as performance metrics, where MD evaluates the precision of drag editing, and IF measures the consistency between the generated and original images by averaging the learned perceptual image patch similarity (LPIPS) [31]. Specifically, 1-LPIPS is employed as the IF metric in our experiment to facilitate comparison. In addition, we compare the average time required per point to demonstrate the time efficiency of our proposed FastDrag. The results are given in Fig. 8.
@@
Apart from [28],[16],[15], two other state-of-the-art methods, i.e., GoodDrag [32] and DiffEditor [19], are also adopted for comparison, with DiffEditor being the current fastest drag-based editing method. Due to well-designed one-step warpage optimization and consistency-preserving strategy, our FastDrag does not require LoRA training preparation, resulting in significantly reduced time consumption (i.e., 3.12 seconds), which is nearly 700% faster than DiffEditor (i.e., 21.68 seconds), and 2800% faster than the typical baseline DragDiffusion (i.e., 1 min and 21.54 seconds). Moreover, even using standard U-Net without LCM, our method is still much faster than DiffEditor and far outperforms all other state-of-the-art methods. It is particularly noteworthy that, even with an A100 GPU, DiffEditor still requires 13.88 seconds according to [19], whereas FastDrag only requires 3.12 seconds on an RTX 3090.
@@ 4.3 Ablation Study
Inversion Step: To determine the number of inversion steps in diffusion inversion with LCM-equipped U-Net, we conduct an ablation experiment with number of inversion steps set as $t$ = 4, 6, 8, 10, 12, 14, 20, and 30, where IF and MD are used to evaluate the balance between the consistency with original image and the desired drag effects. The results are given in Fig. 9 and Fig. 8, where we can see that when $t<6$, the generated images lack sufficient detail to accurately reconstruct the original images. Conversely, when $t>6$, it can successfully recover complex details such as intricate fur textures and dense stone while maintaining high image quality. However, when $t>14$, some image details lost, which negatively impacts the effectiveness of the drag effect. By comprehensive evaluation of both the drag effect and the similarity to the original images, we select 10 as the number of inversion steps for our method with LCM to balance the drag effect.
@@
5 Limitations
Despite FastDrag’s impressive editing speed compared to SOTA methods, it shares some common limitations. 1) Overly Smooth and Finer Details Loss: Similar to other diffusion-based methods [3, 28], FastDrag occasionally loses fine textures from the original images, as shown in Fig. 6, row 4. Despite this, FastDrag outperforms other methods in speed and overall performance. 2) Extremely Long-distance Drag Editing: In such case, object details may be lost due to the lower-dimensional latent space, in which significant changes in detail (i.e., long-drag editing) can disrupt the semantics, making it harder to preserve all details. Nevertheless, FastDrag handles long-distance editing better than other SOTA methods, as illustrated in Fig. 12, where our method successfully achieves long-distance drag editing that others fail to achieve. 3) Highly Relying on Precise Drag Instruction: Achieving optimal results depends heavily on clear drag instructions. As with other SOTA methods, precise input, such as excluding irrelevant areas from the mask (e.g., the face in Fig. 13, row 2) or correctly placing the handle point (e.g., beak in Fig. 13), is essential for better performance.
Figure 12: Illustration of failure cases for limitation analysis under extremely long-distance drag editing. Our FastDrag method may lose some detailed information in these cases but still achieves better editing performance compared to state-of-the-art (SOTA) methods.
Figure 13: Illustration of the limitation analysis with failed and successful drag editing for highly relying on precise drag instruction. (a) It is best to exclude the face from the mask region. (b) The handle point should ideally be placed where the “beak" feature is more prominent.
@@ 6 Conclusion
7 Acknowledgments
This work was partly supported by Beijing Nova Program (20230484261).
@@ Appendix A Supplementary Experiments
Figure 16: Overall editing time comparison with different diffusion steps between FastDrag and DragDiffusion. All experiments are conducted on RTX 3090 with diffusion step set as 1, 20, and 50 respectively. Optimization means latent optimization.
Appendix B Single-step Diffusion
When integrating DragDiffusion with a single-step diffusion model, the editing time is still much longer than that of FastDrag. For DragDiffusion and FastDrag under diffusion steps of 1, 20, and 50, we calculate the time required for inversion, sampling, and latent optimization respectively. The results provided in Fig. 16 show that even with a single diffusion step (i.e., diffusion step set as 1), DragDiffusion still requires significantly more time (20.7 seconds) compared to FastDrag (2.88 seconds).
In addition, as observed in Fig. 16, DragDiffusion spends significantly more time on latent optimization compared to diffusion inversion and sampling. Therefore, reducing the time spent on latent optimization is crucial for minimizing overall editing time, which is precisely what FastDrag accomplishes.
Appendix C Statistical Rigor
To further validate the superiority of our work and to achieve statistical rigor, we conducted an additional experiment by repeating our experiment 10 times under the same experimental settings. We observed that the variances of the performance metrics obtained from 10 realizations of our FastDrag are MD (0.000404), 1-LPIPS (9.44E-11), and Time (0.018), all of which fall within a reasonable range. These statistical results further demonstrate the effectiveness and stability of our method for drag editing.
Appendix D Implementation Details
We utilize a pretrained LDM (i.e., Stable Diffusion 1.5 [27]) as our diffusion model, where the U-Net structure is adapted with LCM-distilled weights from Stable Diffusion 1.5. It is worth emphasizing that the U-Net structure used in our model is widely used in image generation methods [28, 15, 16, 32, 4]. Unless otherwise specified, the default setting for inversion and sampling step is 10. Following DragDiffusion [28], classifier-free guidance (CFG) [8] is not applied in diffusion model, and we optimize the diffusion latent at the 7th step. All other configurations follow that used in DragDiffusion. Our experiments are conducted on an RTX 3090 GPU with 24G memory.
Appendix E More Results
@@
Appendix F Societal Impacts
FastDrag has the potential to bring about several positive societal impacts. Firstly, they offer intuitive and efficient image editing tools, catering to artists, designers, and creators, thereby fostering creativity and innovation. Secondly, their user-friendly nature simplifies the image editing process, increasing accessibility and participation among a wider audience. In addition, our FastDrag improves efficiency and saves the user’s time and effort, thus increasing productivity. Moreover, the innovative and flexible nature of these methods opens up possibilities for various applications, spanning art creation, design, education, and training.
However, along with these positive aspects, FastDrag can also have certain negative societal implications. They may be exploited by unethical individuals or organizations to propagate misinformation and fake imagery, potentially contributing to the spread of false news and undermining societal trust. Furthermore, widespread use of image editing tools may encroach upon individual privacy rights, particularly when unauthorized information or imagery is manipulated. Moreover, inappropriate or irresponsible image editing practices could lead to social injustices and imbalances, such as distorting facts or misleading the public, thereby influencing public opinion and policies negatively.